# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v1

@@ Abstract
We present both a theoretical and a methodological framework that addresses a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities, such as the electronic-structure Hamiltonian. Inspired by covariant theory in physics, we address this problem by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this foundation, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove. Our approach offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness in deep learning methodologies. We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, demonstrating state-of-the-art performance across six benchmark databases.
@@
Deep Learning Theory, Neural Networks, SO(3)-equivariance, Science for AI, AI for Science
@@ 1 Introduction
With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research [1]. For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods [2, 3] suffer from computational complexities of $\mathcal{O}(N^{3})$. In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ [4, 5, 6, 7], accelerating processes by thousands or even millions of times. This reduction in computational demand not only unlocks possibilities for analyzing extremely large atomic systems but also paves the way for efficient material simulation, discovery, and design that were previously unimaginable.
@@
However, deep learning methods still face substantial challenges when processing physical systems. To align with fundamental physical laws, these methods must strictly adhere to symmetry principles. For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group. Traditional deep learning methods [8, 9] typically approximated SO(3)-equivariance through rotational data augmentation. However, this approach does not provide the strict assurance of SO(3)-equivariance required for physical research. Besides, the calculation of physical quantities calls for ultra-high numerical accuracy. For example, it is typically required that the errors on Hamiltonian calculation being restricted within the sub-millielectron volts (sub-meV) scale, to ensure accurate derivation of downstream physical quantities and related material properties. This necessitates that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets. However, existing deep learning methods are unable to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) leads to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. Although several efforts have been made to address this challenge—introducing gated activation mechanisms [7, 10, 11, 12], employing discrete spherical harmonic bases [13, 14, 15], and developing hybrid regression approaches [16]—they still fall short, to varying degrees, of fully reconciling strict adherence to SO(3)-equivariance with the requirements for high-precision generalization capability. As analyzed in previous literature [13], this issue is commonly found in machine learning tasks that demand both strict equivariance and fine-grained generalization performance.
@@
To address this challenge, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning. We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties. For example, in special relativity [17], the spacetime interval between events is an invariant under Lorentz transformations. This invariance is foundational in formulating physical laws that involve covariant quantities. Furthermore, from the perspective of deep representation learning, the attribute of preserving invariance under non-linear operations is a significant advantage, due to its compatibility with non-linear expressive capabilities. Built upon these insights, we theoretically solve the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations from these SO(3)-invariant ones. Specifically:
@@
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations as well as the regression target $\mathbf{Q}$, from high-quality SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove.
We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic structure Hamiltonians, which are central to quantum physics and inherently complex, presenting substantial challenges for regression models. Our method significantly enhances the state-of-the-art performance on six benchmark Hamiltonian databases, demonstrating that our approach not only maintains strict SO(3)-equivariance but also unleashes powerful non-linear expressiveness, thereby boosting the prediction accuracy of Hamiltonians, providing an effective deep learning tool for efficient and accurate electronic structure calculations of materials.
Our theory and method for bridging symmetry and expressiveness of deep learning are general and can be easily applied on predicting other physical quantities such as forces, which also obey SO(3)-equivariance. Furthermore, it can also be extended to other application areas beyond physics. For instance, in some computer vision tasks, such as autonomous driving [18], where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation. However, this does not guarantee absolute safety. Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks.
2 The Equivariance-Expressiveness Dilemma in Existing Deep Learning Methods
In this section, we introduce existing efforts to harmonize the deep learning paradigm with SO(3)-equivariance, analyzing the dilemma between strict SO(3)-equivariance and non-linear expressive capabilities.
2.1 Approaches with Strict SO(3)-equivariance but Weak Expressiveness
Most existing methods have focused on achieving SO(3)-equivariance at the neural mechanism level by applying group theory-guaranteed SO(3)-equivariant operators [19, 20]. These operators, including linear scaling, element-wise sum, direct products, direct sums, and Clebsch-Gordan decomposition, have been used to build graph neural network architectures for tasks in 3D point cloud analysis [21], molecular dynamic simulation and property prediction [11, 10], as well as Hamiltonian prediction [22, 4, 7, 12]. However, these operators could not provide sufficient expressiveness to fit complex non-linear mappings. The majority of these operators are linear operations, and the only one (direct product) that introduces some non-linearity can only approximate a simple squaring function, and cannot be extended to more complex non-linear forms due to the huge computational burden and difficulty in convergence. Even worse, directly applying highly-expressive non-linear activation functions, such as $Sigmoid$, $Softmax$, and $SiLU$, to SO(3)-equivariant features may result in the loss of equivariance. This dilemma leads to an unsatisfactory trade-off between strict equivariance and non-linear expressiveness, ultimately reducing the achievable performance of these regression models. To alleviate this issue, these methods introduced a gated activation mechanism that processes SO(3)-invariant features ($l=0$) using non-linear activations and employs them as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance. However, from the perspective of SO(3)-equivariant features, this mechanism is a combination of linear operations and cannot fundamentally increase their non-linear expressive capabilities.
2.2 Approaches with Remarkable Expressiveness but Approximate SO(3)-equivariance
Li et al. [6] proposed a local coordinate strategy to achieve SO(3)-equivariance. This strategy relied on the transformation relationship between local and global coordinates to achieve equivariance, without imposing restrictions on the neural network, allowing the use of network modules that possess non-linear expressiveness. However, this strategy, solely depending on coordinate transformations, only works under holistic rigid rotations of the atomic system. Without exploration of a neural mechanism for equivariance, when the local coordinates experience non-rigid perturbations such as thermal movements or bilayer twists, this strategy offers no guarantees on maintaining symmetry properties; Zitnick et al. [13] decomposed features into weighted sums of spherical harmonic basis functions, each corresponding to a discrete point on a unit sphere. Such discrete decomposition enables non-linear operations on the coefficients to enhance expressiveness. In subsequent developments [14, 15, 23], this approach has demonstrated a remarkable capacity to fit complex functions. However, it degenerates from continuous to discrete rotational equivariance, no longer maintaining strict equivariance to continuous rotational operations; Yin et al. [16] proposed a hybrid framework consisting of both group theory-guaranteed SO(3)-equivariant mechanisms and non-linear mechanisms to regress Hamiltonians. In this framework, the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy. However, the equivariance achieved through data-driven methods does not have strict theoretical guarantees. In summary, although these methods with approximate SO(3)-equivariance achieve high numerical accuracy due to their extensive use of non-linear mechanisms, they are best suited for the applications that primarily require numerical approximations but not for tasks that demand theoretical depth in physics. In some applications, the demands for symmetry are extremely high. Even minor deviations from perfect symmetry can result in incorrect physical results.
Combining the analysis in this section, existing deep learning methods have compromised either strict SO(3)-equivariance or non-linear expressiveness when processing physical systems. In this paper, we develop a novel framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks. This solution resolves the equivariance-expressiveness dilemma by deeply exploring the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities.
@@ 3 Problem Formalization
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear mechanisms, such as $Sigmoid$, $Softmax$ and $Silu$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, is the core problem this paper aims at solving.
@@ 4 Theory
This section establishes a rigorous theoretical foundation for integrating SO(3)-equivariance with non-linear expressiveness in the deep learning paradigm. First, we propose a theoretical construct to derive SO(3)-invariant quantities from the SO(3)-equivariant targets, aiming at facilitating the learning of high-quality SO(3)-invariant features. Second, we propose a gradient-based mechanism encoding features maintaining theoretically SO(3)-equivariance while possessing the non-linear expressive capability induced from SO(3)-invariant features to capture the complex non-linear patterns inherent in physical quantities.

### v3

@@ Abstract
We present both a theoretical and a methodological framework that addresses a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities. Inspired by covariant theory in physics, we address this problem by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this foundation, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove. We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of elements and challenging scenarios show dramatic breakthroughs on the state-of-the-art prediction accuracy, with improvements of up to 40% in predicting Hamiltonians and up to 76% in predicting downstream physical quantities such as occupied orbital energy. Our approach goes beyond handling physical systems and offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm.
@@
Deep Learning, Neural Networks, SO(3)-equivariance, Atomic Modeling, Hamiltonian Prediction
@@ 1 Introduction
With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research [1]. For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods [2, 3] suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system. In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ [4, 5, 6, 7, 8], accelerating processes by thousands or even millions of times. This reduction in computational demand not only unlocks possibilities for analyzing extremely large atomic systems but also paves the way for efficient material simulation, discovery, and design that were previously unimaginable.
@@
However, deep learning methods still face substantial challenges when processing physical systems. To align with fundamental physical laws, these methods must strictly adhere to symmetry principles. For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group. Traditional deep learning methods [9, 10] typically approximated SO(3)-equivariance through rotational data augmentation. However, this approach does not provide the strict assurance of SO(3)-equivariance required for physical research. Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets. However, existing deep learning methods are unable to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. As pointed out in previous literature [11], this issue is commonly found in machine learning tasks that demand both strict equivariance and fine-grained generalization performance. Although several efforts have been made to alleviate this issue by introducing gated activation mechanisms [8, 7, 12, 13, 14], developing local coordinate strategy [6], and employing discrete spherical channels [11, 15, 16], they still, to varying degrees, compromised either non-linear expressiveness [7, 8, 12, 13, 14] or strict SO(3)-equivariance [11, 15, 16] when processing physical systems. Please refer to Supplementary Information A for an in-depth analysis.
@@
To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning. We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties. For example, in special relativity [18], the spacetime interval between events is an invariant under Lorentz transformations. This invariance is foundational in formulating physical laws that involve covariant quantities. Furthermore, from the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities. Built upon these insights, we theoretically solve the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations from these SO(3)-invariant ones. Specifically:
@@
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$, from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove.
We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deeply understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as shown in Fig. 1. Our method dramatically enhances the state-of-the-art performance on eight benchmark Hamiltonian databases from the renowned DeepH and QH9 benchmark series [6, 8, 19], showing excellent generality to various challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories, with an accuracy improvement of Hamiltonians of up to 40%. Moreover, our approach substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonians including occupied orbital energies and electronic wavefunctions, with an accuracy improvement by a maximum of 76% and 17%, respectively. Our leading performance comprehensively demonstrate that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
Figure 1: Illustration of the high-dimensional complexity of the Hamiltonian. In an atomic system, a Hamiltonian matrix is defined between each pair of atoms within each other’s cutoff radius, constituting a large cluster of Hamiltonian matrices for all atom pairs. Besides, each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 2 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here. Thus, for a Hamiltonian prediction model, the space of the regression target is typically huge.
@@ 2 Problem Formalization
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $Silu$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, is the core problem this paper aims at solving.
@@ 3 Theory
This section establishes a rigorous theoretical foundation for integrating SO(3)-equivariance with non-linear expressiveness in the deep learning paradigm. First, we propose a theoretical construct to derive SO(3)-invariant quantities from the SO(3)-equivariant targets, aiming at facilitating the learning of high-quality SO(3)-invariant features. Second, we propose a gradient-based mechanism encoding features maintaining theoretically SO(3)-equivariance while possessing the non-linear expressive capability induced from SO(3)-invariant non-linear features to capture the complex non-linear patterns inherent in physical quantities.